# GETS: Ensemble Temperature Scaling for Calibration in Graph Neural Networks

### v1

@@ 4.2 Ensemble Calibration by Graph MoE
The node-wise ensemble idea mentioned in the last section coincides with the so called MoE framework (Jordan & Jacobs, 1994; Jacobs et al., 1991; Jacobs, 1995; Wang et al., 2024). MoE allows for specialization by assigning different experts to handle distinct inputs, enabling each expert to focus on specific factors. This specialization is particularly beneficial when inputs are diverse, as it allows experts to learn the unique properties of each factor more effectively than one single GAT model (Hsu et al., 2022) for many inputs at the same time. Moreover, MoE employs sparse activation through a gating network, which selectively activates only relevant experts for each input, reducing computational overhead compared to the dense computation in GATs (Shazeer et al., 2017; Lepikhin et al., 2020). The overall framework design is shown in Figure 2 (Zhang et al., 2020).
To adapt MoE for graph-based calibration tasks, we propose the GETS framework that integrates both the input ensemble and the model ensembles. In GNNs like GCNs, nodes update their representations by aggregating information from their neighbors. The standard GCN propagation rule is (Wang et al., 2024):
@@
where $\mathbf{h}_{i}^{\prime}$ is the updated feature vector for node $i$ after the GCN layer, $\mathcal{N}_{i}$ is the set of neighboring nodes of node $i$. We train different GCNs based on various ensemble of inputs as different experts, $\mathbf{h}_{j}$ is the feature vector of the neighboring node $j$ and $\mathbf{W}$ is a learnable matrix for the GCN network. $\frac{1}{\sqrt{|N_{i}||N_{j}|}}$ is the normalization term where $|N_{i}|$ and $|N_{j}|$ are the degrees (number of neighbors) of nodes $i$ and $j$, respectively. In many GNN papers, the normalization term is also denoted as $\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}$ with $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix with added self-loops. $\tilde{\mathbf{D}}$ is the corresponding degree matrix with $\tilde{\mathbf{D}}_{ii}=\sum_{j}\tilde{\mathbf{A}}_{ij}$ and $\sigma$ is a non-linear activation function.
@@
where $\mathbf{W}_{g}$ and $\mathbf{W}_{n}$ are learnable weight matrices control clean and noisy scores, respectively. Noisy gating $\epsilon\sim\mathcal{N}(0,1)$ introduces Gaussian noise for exploration. In the context of calibration, we adapt GETS to combine multiple calibration experts, each specialized in handling specific calibration factors. For those unselected models the gating function $G(\cdot)$ sets them to zeroes to be consistent with Equation 9, for example, if we select the best 2 experts, the rest $M-2$ outputs, the gating function $G(\cdot)$ assigns them to 0. Finally, for each node $i$, the calibrated logits $z_{i}^{\prime}$ are computed as, in a similar form as Equation 5:
@@
Each calibration expert $g_{m}$ can implement different node-wise calibration strategies, where node-wise temperatures are usually included in the model. The model is trained by minimizing a calibration loss over all nodes: $L=\frac{1}{N}\sum_{i=1}^{N}\ell_{\text{cal}}\left(z_{i}^{\prime},y_{i}\right)$, where $\ell_{\text{cal}}$ is a differentiable calibration loss function (e.g., cross-entropy loss) mentioned before and $N$ is the number of nodes. By leveraging GETS, we enable the calibration model to specialize in handling diverse calibration factors via different experts and adaptively select the most relevant experts for each node through the gating mechanism using node features and topology.
@@ 5.1 Experimental Setup
We include the 10 commonly used graph classification networks for thorough evaluation, The data summary is given in Table 1, refer to Appendix A.2 for their sources. The train-val-test split is 20-10-70 (Hsu et al., 2022; Tang et al., 2024), note that uncertainty calibration models are trained on validation set, which is also referred as the calibration set. We randomly generate 10 different splits of training, validation, and testing inputs and run the models 10 times on different splits.
@@ 5.4 Expert Selection
Figure 3: Illustration of computational efficiency and expert selection properties. (a): Elepsed time for training each model for 10 runs; (b): Primary and secondary expert selections across datasets for various input ensembles. The top bar plot shows the frequency of expert selection, highlighting the significance of combining logits and features in calibration across datasets.
@@
The results show a diverse range of expert choices, with the combination of logits and features $[z,x]$ frequently chosen as the primary experts, particularly for complex datasets like Ogbn-arxiv and Reddit. In contrast, degree-based experts ($d$) are often selected as secondary. For smaller datasets such as Cora and Pubmed, individual expert selections are more prevalent. Noted that only using logits $z$ as the inputs are not preferred by the gating function, which indicates the importance of ensembling different inputs. These findings indicate that integrating multiple input types—especially logits and features—improves calibration, with the optimal expert combination depending on the complexity of the dataset.
@@ 5.5 Ablation Studies
Generally, choosing GAT and GIN as the models to train experts does not provide significant advantages over using GCN, as shown in Table 3. GETS-GAT yields comparable performance to GATS, also encounters the same OOM issue in the Reddit dataset, demonstrating its limitations in handling large-scale datasets. Furthermore, while GETS-GIN shows improved results on datasets like Cora-full, it underperforms on several other datasets, such as Citeseer, Computers, and Reddit, compared to GCN-based calibration. Notably, in simpler datasets or those with smaller graph structures, GCN tends to achieve better calibration results without overfitting. This suggests that a simpler GNN-based calibrator, like GCN, is preferable for training experts as it offers a more balanced trade-off between model complexity and calibration performance, while avoiding potential issues like overfitting, which may arise when more complex GNN architectures are used.
@@ A.1.1 Base GNN Classifier Settings
During training, we use a learning rate of $1\times 10^{-2}$. We tune the weight decay parameter to prevent overfitting and consider adding early stopping with a patience of 50 epochs. The model is trained for a maximum of 200 epochs to ensure convergence. The specifics are summarized in Table 4.
@@ A.1.2 GETS Parameter Settings
GETS includes an expert selection mechanism that incorporates logits, features, and node degree information, either individually or in combination, depending on the configuration that best fits the dataset. For all experiments, we maintain a fixed bin size of 10 for uncertainty estimation. The calibrators are trained with a learning rate of either $1\times 10^{-1}$ or $1\times 10^{-3}$, depending on the model’s complexity, and we tune weight decay ranging from 00 to $0.1$ to further regularize the training process. The calibrators are trained for up to 1000 epochs, with early stopping applied based on a patience of 50 epochs.
@@ A.1.3 Baseline Models Settings
TS (Guo et al., 2017) adjusts the logits output by a model using a single scalar temperature parameter $T>0$ The calibrated logits $z^{\prime}_{i}$ for sample $i$ are computed as:
@@ A.3 Algorithmic fairness among degree groups
A.4 Ablatiion of Ensemble Strategies
@@ A.5 Calibration Reliability Diagram Results
Based on classifier GCN, the reliability diagrams reflect how the confidence aligns with the accuracies in each bin. Reliability diagram of calibration results of different models are given in Figure 4 and 5. The ECE calculated by degree based binning is shown in Figure 6.

### v2

@@ 4.2 Ensemble Calibration by Graph MoE
The node-wise ensemble idea mentioned in the last section coincides with the so-called MoE framework (Jordan & Jacobs, 1994; Jacobs et al., 1991; Jacobs, 1995; Wang et al., 2024). MoE allows for specialization by assigning different experts to handle distinct inputs, enabling each expert to focus on specific factors. This specialization is particularly beneficial when inputs are diverse, as it allows experts to learn the unique properties of each factor more effectively than one single GAT model (Hsu et al., 2022) for many inputs at the same time. Moreover, MoE employs sparse activation through a gating network, which selectively activates only relevant experts for each input, reducing computational overhead compared to the dense computation in GATs (Shazeer et al., 2017; Lepikhin et al., 2020). The overall framework design is shown in Figure 2 (Zhang et al., 2020). To adapt MoE for graph-based calibration tasks, we propose the GETS framework that integrates both the input ensemble and the model ensembles. In GNNs like GCNs, nodes update their representations by aggregating information from their neighbors. The standard GCN propagation rule is (Wang et al., 2024):
@@
where $\mathbf{h}_{i}^{\prime}$ is the updated feature vector for node $i$ after the GCN layer, $d_{i}$ is the set of neighboring nodes of node $i$. We train different GCNs based on various ensemble of inputs as different experts, $\mathbf{d}_{j}$ is the feature vector of the neighboring node $j$, and $\mathbf{W}$ is a learnable matrix for the GCN network. $\frac{1}{\sqrt{|d_{i}||d_{j}|}}$ is the normalization term where $|d_{i}|$ and $|d_{j}|$ are the degrees (number of neighbors) of nodes $i$ and $j$, respectively. In many GNN papers, the normalization term is also denoted by $\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}$ with $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix with added self-loops. $\tilde{\mathbf{D}}$ is the corresponding degree matrix with $\tilde{\mathbf{D}}_{ii}=\sum_{j}\tilde{\mathbf{A}}_{ij}$ and $\sigma$ is a non-linear activation function.
@@
where $\mathbf{W}_{g}$ and $\mathbf{W}_{n}$ are learnable weight matrices that control clean and noisy scores, respectively. Noisy gating $\epsilon\sim\mathcal{N}(0,1)$ introduces Gaussian noise for exploration. In the context of calibration, we adapt GETS to combine multiple calibration experts, each specialized in handling specific calibration factors. For those unselected models the gating function $G(\cdot)$ sets them to zeroes to be consistent with Equation 9, for example, if we select the best 2 experts, the rest $M-2$ outputs, the gating function $G(\cdot)$ assigns them to 0. Finally, for each node $i$, the calibrated logits $z_{i}^{\prime}$ are computed as, in a similar form as Equation 5:
@@
Each calibration expert $g_{m}$ can implement different node-wise calibration strategies, where the model usually includes node-wise temperatures. The model is trained by minimizing a calibration loss over all nodes: $L=\frac{1}{N}\sum_{i=1}^{N}\ell_{\text{cal}}\left(z_{i}^{\prime},y_{i}\right)$, where $\ell_{\text{cal}}$ is a differentiable calibration loss function (e.g., cross-entropy loss) mentioned before and $N$ is the number of nodes. By leveraging GETS, we enable the calibration model to specialize in handling diverse calibration factors via different experts and adaptively select the most relevant experts for each node through the gating mechanism using node features and topology.
@@ 5.1 Experimental Setup
We include the 10 commonly used graph classification networks for a thorough evaluation, The data summary is given in Table 1, refer to Appendix A.2 for their sources. The train-val-test split is 20-10-70 (Hsu et al., 2022; Tang et al., 2024), note that uncertainty calibration models are trained on the validation set, which is also referred to as the calibration set. We randomly generate 10 different splits of training, validation, and testing inputs and run the models 10 times on different splits.
@@ 5.4 Expert Selection
Figure 3: Illustration of computational efficiency and expert selection properties. (a): Elapsed time for training each model for 10 runs; (b): Primary and secondary expert selections across datasets for various input ensembles. The top bar plot shows the frequency of expert selection, highlighting the significance of combining logits and features in calibration across datasets.
@@
The results show a diverse range of expert choices, with the combination of logits and feature $[z,x]$ frequently chosen as the primary experts, particularly for complex datasets like Ogbn-arxiv and Reddit. In contrast, degree-based experts ($d$) are often selected as secondary. For smaller datasets such as Cora and Pubmed, individual expert selections are more prevalent. Note that only using logits $z$ as the inputs is not preferred by the gating function, which indicates the importance of ensembling different inputs. These findings indicate that integrating multiple input types—especially logits and features—improves calibration, with the optimal expert combination depending on the complexity of the dataset.
@@ 5.5 Ablation Studies
Generally, choosing GAT and GIN as the models to train experts does not provide significant advantages over using GCN, as shown in Table 3. GETS-GAT yields comparable performance to GATS, and also encounters the same OOM issue in the Reddit dataset, demonstrating its limitations in handling large-scale datasets. Furthermore, while GETS-GIN shows improved results on datasets like Cora-full, it underperforms on several other datasets, such as Citeseer, Computers, and Reddit, compared to GCN-based calibration. Notably, in simpler datasets or those with smaller graph structures, GCN tends to achieve better calibration results without overfitting. This suggests that a simpler GNN-base calibrator, like GCN, is preferable for training experts as it offers a more balanced trade-off between model complexity and calibration performance, while avoiding potential issues like overfitting, which may arise when more complex GNN architectures are used.
@@ 6 Conclusion
Acknowledgments
Dingyi Zhuang acknowledges the support by the U.S. Department of Energy’s Office of Energy Efficiency and Renewable Energy (EERE) under the Vehicle Technology Program Award Number DE-EE0009211. Yunhan Zheng acknowledges the support by the National Research Foundation (NRF), Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. The Mens, Manus, and Machina (M3S) is an interdisciplinary research group (IRG) of the Singapore MIT Alliance for Research and Technology (SMART) centre. Shenhao Wang acknowledges the support of Research Opportunity Seed Fund (ROSF2023) from the University of Florida.
@@ A.1.1 Base GNN Classifier Settings
During training, we use a learning rate of $1\times 10^{-2}$. We tune the weight decay parameter to prevent overfitting and consider adding early stopping with patience of 50 epochs. The model is trained for a maximum of 200 epochs to ensure convergence. The specifics are summarized in Table 4.
@@ A.1.2 GETS Parameter Settings
GETS includes an expert selection mechanism that incorporates logits, features, and node degree information, either individually or in combination, depending on the configuration that best fits the dataset. For all experiments, we maintain a fixed bin size of 10 for uncertainty estimation. The calibrators are trained with a learning rate of either $1\times 10^{-1}$ or $1\times 10^{-3}$, depending on the model’s complexity, and we tune weight decay ranging from 00 to $0.1$ to regularize the training process further. The calibrators are trained for up to 1000 epochs, with early stopping applied based on patience of 50 epochs.
@@ A.1.3 Baseline Models Settings
TS (Guo et al., 2017) adjusts the logits output by a model using a single scalar temperature parameter $T>0$. The calibrated logits $z^{\prime}_{i}$ for sample $i$ are computed as:
@@ A.3 Algorithmic fairness among degree groups
A.4 Ablation of Ensemble Strategies
@@ A.5 Calibration Reliability Diagram Results
Based on the classifier GCN, the reliability diagrams reflect how the confidence aligns with the accuracies in each bin. The reliability diagram of calibration results of different models is given in Figure 4 and 5. The ECE calculated by degree-based binning is shown in Figure 6.
@@
A.6 GPU Memory Usage
We summarize the GPU memory reserved by each of the calibration methods. By default, all the calibration methods run on the GCN classifier outputs. The memory usage results are shown in Table 8.
Table 8: GPU reserved memory usage for each algorithm. The OOM case is marked with the attempted reserved memory from the algorithm.
Dataset Classifier TS VS ETS CaGCN GATS GETS
Citeseer GCN 164 MB 164 MB 118 MB 164 MB 168 MB 184 MB
Computers GCN 210 MB 210 MB 190 MB 212 MB 348 MB 250 MB
Cora GCN 98 MB 98 MB 98 MB 98 MB 102 MB 110 MB
Cora-full GCN 1380 MB 1380 MB 1380 MB 1380 MB 1384 MB 2018 MB
CS GCN 1546 MB 1546 MB 620 MB 1060 MB 1066 MB 1546 MB
Ogbn-arxiv GCN 3382 MB 3382 MB 3382 MB 1328 MB 2858 MB 2418 MB
Photo GCN 186 MB 186 MB 130 MB 154 MB 216 MB 186 MB
Physics GCN 2368 MB 2368 MB 2368 MB 2368 MB 2390 MB 3504 MB
Pubmed GCN 224 MB 224 MB 224 MB 166 MB 176 MB 224 MB
Reddit GCN 7208 MB 7208 MB 7208 MB 7206 MB 17.54 GB 7208 MB
A.7 SimCalib Comparison
We further compare with the work of SimCalib (Tang et al., 2024), as shown in Table 9.
Table 9: ECE Comparison of GETS and SimCalib results for various datasets and classifiers. Smaller values are highlighted in bold.
Dataset Classifier GETS SimCalib
Cora GCN 2.29 $\pm$ 0.52 3.32 $\pm$ 0.99
GAT 2.00 $\pm$ 0.48 2.90 $\pm$ 0.87
Citeseer GCN 2.50 $\pm$ 1.42 3.94 $\pm$ 1.12
GAT 1.98 $\pm$ 0.30 3.95 $\pm$ 1.30
Pubmed GCN 1.90 $\pm$ 0.40 0.93 $\pm$ 0.32
GAT 0.78 $\pm$ 0.15 0.95 $\pm$ 0.35
Computers GCN 2.03 $\pm$ 0.35 1.37 $\pm$ 0.33
GAT 1.77 $\pm$ 0.27 1.08 $\pm$ 0.33
Photo GCN 2.07 $\pm$ 0.31 1.36 $\pm$ 0.59
GAT 1.10 $\pm$ 0.25 1.29 $\pm$ 0.55
CS GCN 1.34 $\pm$ 0.10 0.81 $\pm$ 0.30
GAT 1.05 $\pm$ 0.27 0.83 $\pm$ 0.32
Physics GCN 0.87 $\pm$ 0.09 0.39 $\pm$ 0.14
GAT 0.29 $\pm$ 0.05 0.40 $\pm$ 0.13
CoraFull GCN 3.32 $\pm$ 1.24 3.22 $\pm$ 0.74
GAT 1.52 $\pm$ 2.27 3.40 $\pm$ 0.91
A.8 Gating Scores Visualization
We also visualize the stack plot of the gating score changes during training. This will reflect how different experts are preferred during different stages of training. By default, we look into the GETS-GCN case run on the GCN classifier.
(a) Citeseer
(b) Citeseer
(c) Cora
(d) Cora-full
(e) CS
(f) Ogbn-arxiv
(g) Photo
(h) Physics
(i) Pubmed
(j) Reddit
Figure 7: GETS gating scores changes during training.
A.9 Ablation on the input ensemble
A.9.1 Different structural information
We conducted a test to replace degree embedding with betweenness centrality embedding, page rank centrality, clustering coefficient, harmonic centrality, Katz centrality, and Node2Vec embedding. We conduct experiments on all datasets, but some of the datasets are too large to compute the network statistics. We left those datasets empty due to limited time in rebuttal periods.
The results are shown in Table 10. For naming convenience, we use GETS-centrality and GETS-Node2Vec to represent different structural information. The original GETS is also repeated here for comparison. By default, we use the classifier GCN. For Node2Vec embedding, we set by default walk_length=20, num_walks=10, workers=4
Table 10: Ablation studies on different structural information, measured by ECE (cleaned values).
GETS 2.50 $\pm$ 1.42 2.03 $\pm$ 0.35 2.29 $\pm$ 0.52 3.32 $\pm$ 1.24 1.34 $\pm$ 0.10 1.85 $\pm$ 0.22 2.07 $\pm$ 0.31 0.87 $\pm$ 0.09 1.90 $\pm$ 0.40 1.49 $\pm$ 0.07
GETS-Betweenness 7.63 $\pm$ 1.35 5.01 $\pm$ 3.49 3.29 $\pm$ 0.75 3.63 $\pm$ 0.68 2.21 $\pm$ 0.90 / 1.84 $\pm$ 0.17 / 2.18 $\pm$ 0.35 /
GETS-Clustering 7.56 $\pm$ 0.88 3.54 $\pm$ 2.30 3.48 $\pm$ 0.41 3.61 $\pm$ 0.77 1.80 $\pm$ 0.08 / 2.00 $\pm$ 0.46 1.15 $\pm$ 0.16 2.18 $\pm$ 0.30 /
GETS-PageRank 7.72 $\pm$ 1.27 3.88 $\pm$ 1.38 3.25 $\pm$ 0.80 3.55 $\pm$ 0.95 1.83 $\pm$ 0.11 / 1.73 $\pm$ 0.12 1.12 $\pm$ 0.12 2.43 $\pm$ 0.37 /
GETS-Node2Vec 4.25 $\pm$ 1.28 2.99 $\pm$ 0.97 3.06 $\pm$ 0.54 3.88 $\pm$ 1.06 1.82 $\pm$ 0.09 / 1.85 $\pm$ 0.40 1.13 $\pm$ 0.15 2.30 $\pm$ 0.33 /
A.9.2 Different input types
We further ablate on different input types. We ablate one of the three input types and create the input ensemble based on the other two. For example, we ablate $z$ and then construct the input ensemble as $\{\mathbf{x},d,[\mathbf{x},d]\}$. For the convenience of naming, we use GETS-DX, GETS-DZ, and GETS-XZ to represent the input ensemble without including $z$, $\mathbf{x}$, and $d$. The results can be found in Table 11. Generally, incorporating more information in the input types would make the results better.
Table 11: Ablation studies on different input ensembles, measured by ECE (cleaned values).
GETS 2.50 $\pm$ 1.42 2.03 $\pm$ 0.35 2.29 $\pm$ 0.52 3.32 $\pm$ 1.24 1.34 $\pm$ 0.10 1.85 $\pm$ 0.22 2.07 $\pm$ 0.31 0.87 $\pm$ 0.09 1.90 $\pm$ 0.40 1.49 $\pm$ 0.07
GETS-DX 4.00 $\pm$ 0.94 2.76 $\pm$ 0.40 3.27 $\pm$ 0.62 3.60 $\pm$ 0.54 1.86 $\pm$ 0.21 2.15 $\pm$ 0.46 2.12 $\pm$ 0.42 1.06 $\pm$ 0.08 2.02 $\pm$ 0.35 1.55 $\pm$ 0.25
GETS-DZ 4.29 $\pm$ 1.24 3.53 $\pm$ 2.36 2.66 $\pm$ 0.71 3.34 $\pm$ 0.34 1.83 $\pm$ 0.15 2.32 $\pm$ 0.27 1.76 $\pm$ 0.58 0.98 $\pm$ 0.12 2.43 $\pm$ 0.56 1.80 $\pm$ 0.16
GETS-XZ 2.93 $\pm$ 0.72 3.49 $\pm$ 1.47 3.09 $\pm$ 0.47 3.26 $\pm$ 0.40 1.94 $\pm$ 0.52 2.24 $\pm$ 0.19 1.33 $\pm$ 0.15 1.02 $\pm$ 0.07 2.56 $\pm$ 0.55 2.28 $\pm$ 0.84
A.10 Ablation on the experts
We extend the Table 3 by including Multi-layer Perceptron (MLP) as the backbone model with the same layer number for training the algorithm. MLP is structure-unaware, which can serve as the baseline. The MLP results show that the selection of the backbone methods for the experts does not necessarily require graph-based models, which offer the opportunity to include a broader range of models for the graph calibration task.
GETS 2.50 $\pm$ 1.42 2.03 $\pm$ 0.35 2.29 $\pm$ 0.52 3.32 $\pm$ 1.24 1.34 $\pm$ 0.10 1.85 $\pm$ 0.22 2.07 $\pm$ 0.31 0.87 $\pm$ 0.09 1.90 $\pm$ 0.40 1.49 $\pm$ 0.07
GETS-MLP 4.82 $\pm$ 0.85 4.09 $\pm$ 1.51 3.19 $\pm$ 0.65 3.51 $\pm$ 1.02 1.89 $\pm$ 0.10 2.38 $\pm$ 0.17 1.38 $\pm$ 0.46 0.99 $\pm$ 0.15 2.18 $\pm$ 0.33 1.96 $\pm$ 0.48